1.      In a hospital unit, there are 8 nurses and 5 physicians; 7 nurses and 3 physicians are females. In a staff person is selected, find the probability that the subject is nurse or male.

 

Staff                     females               males                   total

Nurse                        7                          1                         8

Physicians                3                         2                         5

                                  10                          3                        13


1
Expert's answer
2022-06-15T05:49:50-0400

"A=\{\text{The selected subject is nurse}\}"

"B=\{\text{The selected subject is male}\}"

"AB=\{\text{The selected subject is nurse and male}\}"

"A\cup B=\{\text{The selected subject is nurse or male}\}"


"|\Omega|=13", "|A|=8", "|B|=3", "|AB|=1".


"\mathbb{P}(A)=\frac{|A|}{|\Omega|}=\frac{8}{13}"

"\mathbb{P}(B)=\frac{|B|}{|\Omega|}=\frac{3}{13}"

"\mathbb{P}(AB)=\frac{|AB|}{|\Omega|}=\frac{1}{13}"


By the probability addition theorem "\mathbb{P}(A\cup B)=\mathbb{P}(A)+\mathbb{P}(B)-\mathbb{P}(AB)".


"\mathbb{P}(A\cup B)=\frac{8}{13}+\frac{3}{13}-\frac{1}{13}=\frac{8+3-1}{13}=\frac{10}{13}"